Find the two consective positive integers such that the sum of their
squares is 85.
I didn't find anything about integers in the section 10-3. I don' understand
I hope someone can explain this to me.
Thanks
Cheryl

Answer by Nate(3500)

integer = i
consecutive integer = i + 1
i^2 + (i + 1)^2 = 85
i^2 + i^2 + 2i + 1 = 85
2i^2 + 2i - 84 = 0
i^2 + i - 42 = 0
(i + 7)(i - 6) = 0
i = -7
and
i = 6
Integers: 6 and 7